### v2

# L-MARS: Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search

Abstract
We present L-MARS (Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search), a system that reduces hallucination and uncertainty in legal question answering through coordinated multi-agent reasoning and retrieval. Unlike single-pass RAG, L-MARS decomposes queries into subproblems, issues targeted searches across heterogeneous sources (Serper web, local RAG, CourtListener case law), and employs a Judge Agent to verify sufficiency, jurisdiction, and temporal validity before answer synthesis. This iterative reasoning–search–verification loop maintains coherence, filters noisy evidence, and grounds answers in authoritative law. We evaluated L-MARS on LegalSearchQA, a new benchmark of 200 up-to-date multiple choice legal questions in 2025. The results show that L-MARS substantially improves the accuracy of the factual data, reduces uncertainty, and achieves higher preference scores for the human and LLM judges. Our work demonstrates that multi-agent reasoning with agentic search is a scalable, reproducible blueprint for deploying LLMs in high-stakes domains requiring precise legal retrieval and deliberation. The code is available at: https://github.com/boqiny/L-MARS.
1 Introduction
Large language models (LLMs) such as GPT (gpt_4o_system_card, ), Claude (anthropic2024claude35sonnet, ), and Gemini (gemini2024arxiv, ) are increasingly applied to legal tasks such as case law retrieval (locke2022caselawretrievalproblems, ), statutory interpretation, and automated legal assistance(rincon2021legalclinic, ). Their ability to generate fluent context-sensitive responses makes them attractive for legal services, but direct application often results in hallucinations (ji2023survey, ; xu2025hallucinationinevitableinnatelimitation, ) or uncertainty (huang2024surveyuncertaintyestimationllms, )—confidently stated yet factually unsupported answers. In legal contexts, such errors carry significant real-world risk, as incorrect citations or outdated statutes can undermine credibility and thus negatively affect decision-making.
Two primary strategies are commonly used to mitigate hallucinations in domain-specific LLMs. Domain-specific fine-tuning(chalkidis2020legalbert, ) can improve performance in legal queries, but is costly and fragile. Legal systems evolve rapidly: new regulations are introduced, existing laws are amended, and obsolete provisions are repealed. Maintaining accuracy would require frequent retraining, which makes this approach impractical for continuously updated domains. Retrieval-Augmented Generation (RAG) (lewis2020rag, ) avoids the retraining burden by retrieving relevant documents from an external corpus at inference time, but its effectiveness depends heavily on retrieval accuracy. If key legal evidence is missed, the model may still hallucinate or produce incomplete reasoning.
To address these challenges, we introduce L-MARS, a system that integrates diverse retrieval sources with structured outputs and multi-turn model reasoning. L-MARS combines online search via the Serper API (serperapi, ) for up-to-date web information, offline retrieval using RAG over a curated local legal database, and case law retrieval through the CourtListener API (courtlistenerapi, ) for authoritative legal opinions. Unlike prior single-turn pipelines, L-MARS employs a multi-agent workflow that iteratively decomposes queries, executes targeted retrieval, and performs sufficiency checks before synthesizing final answers.
Figure 1: Comparison of the two L-MARS operating modes. (a) Simple Mode: executes a single-pass retrieval–summarization pipeline. (b) Multi-Turn Mode: adds Judge Agent–guided iterations with sufficiency checks and query refinement.
L-MARS operates in two modes. In the Simple Mode, a single-turn pipeline retrieves relevant legal documents and uses a model with reasoning capacity to integrate them into the final answer. In the Multi-Turn Mode, a Query Agent refines the user query through clarifying sub-questions, a Judge Agent evaluates the sufficiency and relevance of retrieved evidence, and a Summary Agent synthesizes the final answer. This iterative loop enables the system to adaptively search, verify, and refine until an evidence sufficiency threshold is met.
By combining reasoning with agentic search, L-MARS achieves higher factual accuracy and significantly reduces uncertainty compared to the baseline of the GPT, Claude, and Gemini models. The framework also provides a flexible architecture that can switch between low-latency simple mode and high-accuracy multi-turn mode, offering a reproducible blueprint for high-stakes domains such as law.
Table 1: Examples of Model-Expressed Uncertainty in Legal QA Responses
Model Uncertainty Expressions in Pure LLM Responses for Legal Questions
Regulations may differ across various cities and provinces in China.
As of my last update in October 2023, public protest in China is heavily regulated…
Legal interpretations and implementations may vary between EU Member States.
This information is based on the legal framework as of 2023 and may be subject to change.
It is important to note that regulations are subject to change, especially with new laws or court rulings.
2 Related Work
Agentic RAG and Multi-Agent Framework
RAG grounds the model output in external evidence to reduce hallucination. Moving beyond single-turn pipelines, interleaving reasoning with retrieval (yao2023reactsynergizingreasoningacting, ; khattab2023, ) improves factual grounding and sample efficiency. Multi-agent frameworks (wu2023autogen, ; langgraph2024, ) operationalize role specialization (planner, retriever, verifier) and conversational coordination, enabling iterative retrieval, critique, and revision. Recent agentic RAG variants further combine planning and verification with retrieval to maintain long-horizon reasoning(singh2025agenticRAG, ).
Search Systems for LLMs
Long-horizon tool-using research requires agents that actively plan queries, browse, and synthesize sources. Search-o1 and WebThinker (searcho1, ; li2025webthinkerempoweringlargereasoning, ) integrates agentic search into large reasoning models to trigger retrieval at uncertainty points and reason over documents before injection into the chain. In production settings, OpenAI’s Deep Research (openai2025deepresearch, ) exposes an agentic, web-scale research loop designed for deeper, broader investigations. These systems complement “think-then-answer” reasoning models (openai2024o1arxiv, ; openai2025o3systemcard, ) by coupling deliberation with targeted information acquisition.
Reasoning Models and Techniques
Reasoning-oriented prompting has shown strong gains: chain-of-thought (wei2023chainofthoughtpromptingelicitsreasoning, ) prompting encourages models to externalize intermediate steps, while self-consistency reduces variance by aggregating diverse traces (wang2022selfconsistency, ). Structured interaction methods, including debate-style critique and verifier-based confidence estimation, further improve factuality and calibration (kadavath2022know, ). More recently, large reasoning models such as OpenAI’s o1(openai2024o1arxiv, ), Qwen-QwQ(qwq-32b-preview, ), and DeepSeek-R1(deepseek-r1, ) extend these ideas by scaling the deliberation at the test time, explicitly demonstrating multi-step reasoning traces across domains such as mathematics, code, and science.
Legal AI and Benchmarks
Legal AI has progressed through a series of domain-specific benchmarks and corpora. LegalBench(guha2023legalbench, ) measures practical legal reasoning for diverse tasks ; LexGLUE(chalkidis2022lexglue, ) standardizes the evaluation across multi-task legal NLU; and Pile-of-Law(henderson2022pileoflaw, ) provides a large and responsibly curated pretraining corpus. More recent work has introduced reasoning and retrieval focused evaluations: Reasoning-Focused Legal Retrieval Benchmark(Zheng_2025, ) introduces Bar Exam QA and Housing Statute QA to capture realistic legal RAG settings, while LEXam (fan2025lexambenchmarkinglegalreasoning, ) benchmarks long-form, multi-step legal reasoning with law exam questions in English and German. Our work builds on this trajectory by combining retrieval with agentic multi-turn reasoning and domain-specific verification, complementing these benchmarks with a system architecture explicitly designed to reduce hallucination and uncertainty.
3 Methodology
3.1 Problem Formation
We formalize legal question answering as generating both a reasoning chain $R$ and a final answer $a$ given a user query $q$, external retrieval results $D$, and optional user clarifications $U$. Each query state is represented as $(q,U,D)$, which evolves iteratively through the workflow. At each step, the system decides whether (i) expand $D$ by issuing new search queries, or (ii) terminate with a sufficient answer. Formally, the objective is to learn a mapping
$(q,U,D)\;\mapsto\;(R,a),$
where $R=\{r_{1},r_{2},\dots,r_{T}\}$ is a sequence of reasoning steps interleaved with the retrieved evidence, and $a$ is the final response string. The Judge Agent makes a sufficiency decision $s\in\{\textsc{Sufficient},\textsc{Insufficient}\}$ and refinement notes $E$ guiding the next query state. The loop terminates when $s=\textsc{Sufficient}$ or after $M$ iterations, as illustrated in Figure 1.
3.2 Workflow and Agents Overview
Workflow
L-MARS is a multi-agent workflow implemented in LangGraph with structured output validation and tool use. The system employs a directed acyclic graph (DAG) architecture in which agents are represented as nodes and the control flow is represented as edges with conditional routing. The workflow maintains a centralized WorkflowState that tracks: (i) the evolving query representation, (ii) the accumulated search results with metadata, (iii) the iteration history and refinement notes, and (iv) the intermediate agent outputs. State transitions are deterministic and type-checked, ensuring reproducibility across runs.
Agents
The Query agent parses the user’s question into a structured query result (issue type, key entities, time window, jurisdiction, and initial search intents). The Search Agent executes tool calls to the retrieval backends and returns normalized SearchResult objects (title, url, snippet, optional full content, source type). The Judge Agent performs evidence sufficiency checks and an explicit checklist (factual support, jurisdiction match, temporal specificity, and contradiction analysis). Summary Agent composes the final response with citations and rationale. Detailed agent implementations can be found in the Appendix A.
3.3 Agentic Search
Online Search
We adopt an agentic retrieval policy inspired by Search-o1(searcho1, ), interleaving targeted web queries with stepwise reasoning rather than performing a single problem-oriented fetch. Concretely, the Search Agent triggers retrieval when either (i) the Judge Agent flags a knowledge gap (e.g. missing statute/case authority, date or jurisdiction ambiguity), or (ii) the query classifier predicts that external evidence will reduce answer uncertainty. Once activated, the agent executes one of two complementary search modes against Google Serper, as shown in Table 2 :
Basic Search (titles/snippets) parses organic results into a normalized schema (title, URL, site, date, snippet). This mode is fast and used for query exploration, de-duplication, and document selection under a strict token budget.
Enhanced Search (content extraction) uses the top-$m$ URLs selected by the agent (default m=3m{=}3), then fetches page contents and PDFs with robust headers and timeouts. HTML is cleaned with BeautifulSoup; PDFs are parsed with pdfplumber. To avoid flooding the context, we perform a snippet-anchored extraction: given the search snippet, we locate the best-matching sentence in the full text using a token-level $F_{1}$ overlap and return a fixed window of the surrounding context (2.5k chars). This preserves local evidence while discarding boilerplate and navigation chrome. The extractor also normalizes whitespace, strips punctuation, and truncates long pages to a hard cap to guaranty bounded latency.
Table 2: Comparison of Basic vs. Enhanced (Deep) Search on the query “Can I work remotely in the United States as an F1 student as of 2025?”
Dimension Basic Search (Titles & Snippets) Enhanced Search (Full Content)
Source Types Interstride blog, Reddit, Quora Interstride blog (full text), others truncated
Output Length 1,145 chars, 19 lines 4,154 chars, 15 lines
Example Result “Even if you are working remotely… must have work authorization.” “Remote work for international students on F-1 visas: Any work done on US soil requires authorization…”
Content Depth Snippets only; limited context Full article text (4k cap); snippet-anchored window (2.5k chars)
Latency Fast Slower; scraping + parsing
Use Case Quick exploration, de-duplication Detailed evidence; grounding in authoritative text
Local RAG Indexing
We also maintain a local retrieval index over user-provided documents. Markdown files are segmented into overlapping windows of 500 characters with 100-character stride and indexed using the BM25 ranking function (robertson2009bm25, ). At query time, the retriever returns the top-$k$ segments (default $k=5$), enriched with document titles and section headers to preserve context. The indexing process is dynamic: new files added to the inputs directory are automatically detected and incorporated without requiring a system restart.
CourtListener Integration
To incorporate authoritative case law into the reasoning process, we integrate the CourtListener Legal Search API as a tool, providing access to millions of judicial opinions, oral arguments, and dockets curated by the Free Law Project. When enabled, the Search Agent queries by party name, citation, docket number, or keyword. The API returns structured metadata (case name, citation, court, filing date, opinion text), which we normalize into a uniform schema. This integration allows the system to ground answers in both retrieved secondary sources and primary case law to improve legal credibility.
3.4 L-MARS Operating Mode
Simple Mode
The Simple Mode executes a single-pass pipeline that prioritizes minimal latency. The Query Agent generates structured intents from the user question, which the Search Agent uses to retrieve evidence from enabled sources (Serper web snippets, optional BM25 index, and CourtListener). The Summary Agent then composes an answer with citations in one step.
Multi-Turn Mode
The Multi-Turn mode adds an iterative search–judge–refine loop to improve recall and reduce hallucinations. The Query Agent first proposes clarifying follow-ups; optional user responses are folded into a structured query state. Each iteration generates targeted search queries, runs deep search (top-$3$ with content), and invokes a deterministic Judge Agent (T=0T{=}0) that performs chain-of-thought analysis, source quality checks, and date/jurisdiction/contradiction screening to decide sufficiency. If insufficient, the system refines the query state and repeats up to a fixed maximum. A final Summary Agent composes the answer.
Algorithm 1 L-MARS Multi-Turn Workflow
1:Input: user query $q$, max iterations $M$
[…]
4:Receive user responses $U$;  $Q\leftarrow\textsc{Incorporate}(Q,U)$
[…]
6:for $i=1$ to $M$ do
7:  $S\leftarrow\textsc{GenSearchQueries}(Q,R,E)$
8:  $R_{i}\leftarrow\textsc{DeepSearch}(S,k{=}3,\text{with\_content}{=}\text{True})$
9:  $(\text{suff},\text{notes})\leftarrow\textsc{JudgeAgent}(R\cup R_{i},T{=}0)$ $\triangleright$ CoT, quality, date/jurisdiction, contradiction
10:  $R\leftarrow R\cup R_{i}$;  $E\leftarrow E\cup\text{notes}$
11:  if $\text{suff}=\textsc{Sufficient}$ then
12:   break
13:  else
14:   $Q\leftarrow\textsc{Refine}(Q,\text{notes},U)$
15:  end if
16:end for
[…]
18:return $(a)$
Both modes share the same retrieval backends and dual evaluation stack. The key distinction lies in the graph topology: Simple Mode follows a linear path START $\to$ Query $\to$ Search $\to$ Answer $\to$ END, while Multi-Turn Mode includes a conditional loop with the Judge Agent determining whether to continue retrieval or proceed to answer generation based on evidence sufficiency criteria.
Judge Agent and Stopping Rule
By default, the Judge Agent is instantiated with GPT-o3 for stronger reasoning ability, and runs with temperature =0=0 to reduce variance and improve reproducibility. Its prompt elicits three outputs: (i) a chain-of-thought style rationale, (ii) a structured checklist decision on sufficiency, and (iii) missing-evidence directives that guide the next refinement queries. The loop ends once the decision is made Sufficient or when the maximum iteration limit is reached. In practice, we found the Judge Agent especially useful when search results included unofficial forums (e.g., Reddit, Quora) instead of government domains; in such cases it flagged insufficiency and steered subsequent searches toward more compelling, authoritative sources such as government websites and official regulations.
4 Experiments
4.1 Tasks and Datasets
We introduce LegalSearchQA, a benchmark of 200 legal questions designed to target post-training knowledge gaps. Unlike prior benchmarks such as LegalBench or CUAD(hendrycks2021cuadexpertannotatednlpdataset, ), which evaluate reasoning over pre-provided context, UncertainLegalQA tests the end-to-end ability of a system to retrieve and reason over external legal sources—an essential capability for real-world deployment. To avoid reliance on static pre-training knowledge, all questions explicitly reference the legal status as of 2025, requiring retrieval of up-to-date authorities.
The dataset spans multiple domains, with a strong emphasis on federal executive orders (60%), along with coverage of tax provisions (10%), immigration (10%), federal tax inflation updates (10%), housing reforms (5%) and technology law (5%). This distribution reflects a focus on recent federal regulatory actions while still incorporating broader legislative and policy areas. Jurisdictional coverage includes both US federal law and major state-level developments, emphasizing contexts of regulatory flux, pending rulemaking, and areas where conflicting authorities require active retrieval. The questions are further categorized into requirement checks (48%), permissibility assessments (12%), definitions and scope (10%), policy intent (8%), timeline queries (8%), institutional authority (6%), legal authority references (4%) and eligibility thresholds (4%).
Figure 2: Composition of the LegalSearchQA dataset.
4.2 Evaluation Metrics
We employ a comprehensive evaluation framework that combines automated metrics with LLM-based judgment to capture both quantitative and qualitative aspects of LegalSearchQA.
Accuracy
measures the fraction of multiple choice questions answered correctly against expert-annotated ground truth, reflecting the model’s ability to select the legally correct option.
U-Score
We propose U-Score, a rule-based metric for the evaluation of legal QA systems based on uncertainty. It ranges from 0 to 1 (lower is better) and captures reliability across five complementary dimensions: (i) hedging cues: frequency of uncertainty markers such as “may” or “could”; (ii) temporal vagueness: imprecise or underspecified references to time (e.g., “recently” instead of a year); (iii) citation sufficiency and authority: rewarding the use of primary legal sources (statutes, regulations, case law) over secondary commentary; (iv) jurisdictional specificity: explicit recognition of the relevant legal jurisdiction (e.g., U.S. federal vs. California state law); and (v) decisiveness: whether the answer delivers a clear conclusion rather than remaining equivocal. These components are linearly aggregated with weights:
$\textbf{U-Score}=0.25\,H+0.20\,T+0.25\,(1-C)+0.15\,(1-J)+0.15\,(1-D),$ (1)
where $H,T,C,J,D\in[0,1]$ denote hedging, temporal vagueness, citation sufficiency, jurisdictional specificity, and decisiveness, respectively.
LLM-as-Judge
complements these metrics with qualitative ratings using GPT-o3, evaluating answers along four dimensions: factual accuracy (consistency with current law), evidence grounding (relevance and credibility of cited sources), clarity of reasoning (logical coherence and completeness) and uncertainty calibration (appropriate hedging in genuinely ambiguous areas). Each response is assigned an overall rating of low, moderate, or high, reflecting the holistic quality of legal reasoning. Together, this dual evaluation framework captures both surface-level correctness and deeper reasoning quality.
4.3 Implementation Details
All systems were evaluated on the 200-question LegalSearchQA benchmark using a standardized pipeline. The baseline LLMs (GPT-4o, Claude-4-Sonnet, Gemini-2.5-Flash) were run with temperature =0=0 and a 2048 token limit, using structured output or JSON mode to ensure consistency. Each model produced both a multiple choice answer and a supporting explanation, which were aggregated into structured JSON for evaluation. Performance was assessed along three complementary axes: accuracy, U-Score, and LLM-as-Judge. For reporting, accuracy and U-Score are averaged across all questions, while LLM-as-Judge ratings are determined by majority vote among per-question judgments. The complete evaluation pipeline, including prompts and scripts, is provided in the eval/ directory of our repository.
4.4 Results
Main Results on LegalSearchQA
Table 3 reports aggregate results across three conditions: pure LLM inference, L-MARS in simple mode, and L-MARS in multi-turn mode. Our results highlight three key findings. First, both L-MARS variants substantially outperform pure LLM inference in accuracy (up to 98% vs. 86–89%). Second, the U-Score decreases sharply under L-MARS (from 0.55–0.62 to 0.39–0.42), showing a marked reduction in hedging, vagueness, and unsupported conclusions. Third, LLM-as-Judge ratings consistently rank L-MARS outputs higher, with the multi-turn variant producing more thorough and contextually grounded answers. Finally, we note a trade-off in response time: while baseline LLMs answer within 1-4 seconds, L-MARS incurs higher latency (13.6s for simple, 55.7s for multi-turn), reflecting the added retrieval and reasoning steps. The visualization of the results is shown in Figure 3, and a detailed case study can be found in Appendix B.
Table 3: Average performance of baseline LLMs and our proposed L-MARS framework on the LegalSearchQA benchmark.
Method Accuracy U-Score LLM-as-Judge Response Time (s)
GPT-4o 0.89 0.55 Moderate 1.69
Claude-4-Sonnet 0.88 0.62 Moderate 3.84
Gemini-2.5-Flash 0.86 0.58 Moderate 1.87
L-MARS (simple, ours) 0.96 0.42 High 13.62
L-MARS (multi-turn, ours) 0.98 0.39 High 55.67
4.5 Error Analysis
To better understand the limitations of the model, we conducted a qualitative error analysis of incorrect predictions. We observed two recurring patterns. First, models sometimes default to heuristic reasoning, such as assuming generic timelines or standard legal procedures rather than following the specific statutory text. Second, models occasionally overgeneralized from frequent legal phrases, leading to misclassification when a question involved narrow carving-outs or exceptions. These errors highlight two key challenges: reliance on surface-level priors instead of precise retrieval, and difficulty distinguishing closely related legal categories. Both issues underscore the importance of retrieval grounding and calibrated reasoning in complex legal QA.
Figure 3: Average performance of baseline LLMs and our proposed L-MARS framework on the LegalSearchQA benchmark. Bars report Accuracy (blue) and $1-\text{U-Score}$ (purple), with response times shown alongside the $y$-axis labels. L-MARS substantially improves factual correctness and reduces uncertainty, though with higher latency compared to pure LLM inference.
Human Expert Evaluation
In addition, we conducted blind evaluations with graduate law students on a subset of 50 questions. Human preferences were closely aligned with LLM-as-Judge outcomes, with a 0.92 inter-annotator agreement rate. A recurring failure case was the Judge Agent’s tendency to over-reject partially relevant sources, occasionally triggering unnecessary search iterations. This suggests a direction for future work on trajectory-aware judge models that retain awareness of prior retrieval steps.
5 Conclusion
We introduced L-MARS, a multi-agent workflow that integrates structured reasoning, agentic search, and sufficiency verification for legal question answering. By orchestrating specialized agents in iterative loops of query refinement, targeted retrieval, and evidence-based judgment, L-MARS significantly reduces hallucinations and uncertainty compared to pure LLM baselines. Our evaluation of the LegalSearchQA benchmark demonstrates consistent gains in factual accuracy, decisiveness, and grounding in authoritative sources.
6 Limitations and Future Work
While L-MARS improves reliability, its performance remains bounded by retrieval quality: if search engines or legal databases fail to return relevant authorities, downstream reasoning is impaired. The iterative multi-turn mode also introduces higher latency, which may be prohibitive in real-time legal assistance settings. Moreover, our evaluation focuses on US law; broader cross-jurisdictional and multilingual evaluations are needed to assess generalizability. Finally, the Judge Agent, though effective at flagging insufficiency, can be overly conservative, sometimes triggering unnecessary search iterations.
Appendix
Appendix A Agent Instructions
Query Agent Prompts
# Follow-up Questions Generation Prompt
You are a legal assistant helping users with their legal questions.
User’s original question: {user_query}
Generate 2-3 clarifying follow-up questions that would help you provide better
legal assistance.
Focus on:
- Jurisdiction/location if not specified
- Specific circumstances or context
- Timeline or urgency
- What type of help they need (information, next steps, etc.)
Only ask questions that would significantly help in providing better assistance.
# Search Query Generation Prompt
Based on the user’s legal question and any additional context, generate 2-4 specific
search queries.
User Question: {user_query}
Additional Context: {context}
Generate queries for different sources:
- case_law: For searching legal cases and precedents
- web_search: For general legal information and recent updates
- offline_rag: For searching local legal documents in the inputs folder
Make queries specific and focused to get the most relevant results.
# Query Analysis Prompt
Analyze the user’s legal query and provide a detailed structured breakdown.
User Query: {user_query}
Additional Context: {context}
Analyze and extract:
1. Legal category/area (e.g., contract law, criminal law, family law, employment law)
2. People involved with their roles (plaintiff, defendant, client, witness, etc.)
3. Jurisdiction if mentioned or inferable
4. Urgency level based on the nature of the query
5. Specific legal areas involved
6. Timeline or deadlines if mentioned
7. Additional context about the situation
Be thorough but accurate in your categorization.
Summary Agent Prompt
Create a comprehensive answer to the user’s legal question based on the search results.
User’s Question: {user_query}
Search Results:
{results_content}
Provide:
1. A clear, comprehensive answer
2. Key legal points and considerations
3. Important disclaimers about legal advice
Remember: This is informational only and not legal advice.
Judge Agent Prompt
You are a legal research judge evaluating search results.
This is iteration {iteration_count + 1}.
Original Question: {user_query}
Conversation History:
{conversation_context}
Current Search Results ({len(search_results)} results):
{results_summary}
{prev_evaluations}
EVALUATION PROTOCOL WITH CHAIN-OF-THOUGHT:
1. REASONING (Chain of Thought):
Think step by step about whether the search results answer the user’s specific question.
Consider: What was asked? What information was provided? What is still missing?
2. SOURCE QUALITY CHECK:
Analyze source authority - are there sources from:
- Government (.gov) sites?
- Court decisions/legal databases?
- Educational institutions (.edu)?
- How many authoritative vs user-generated content sources?
3. DATE CHECK:
- Are the sources current and relevant to today’s date?
- If there are older sources, do we also have recent confirmations?
- Flag if critical information might be outdated
4. JURISDICTION CHECK:
- Does the jurisdiction of sources match the user’s location/scope?
- For US federal vs state law, is the distinction clear?
- User mentioned: {conversation_context if conversation_context
else "No specific jurisdiction mentioned"}
5. CONTRADICTION SCAN:
- Do any sources contradict each other?
- If yes, what specific elements conflict?
- Do we need more specific queries to resolve conflicts?
STOP RULE:
Mark as SUFFICIENT when:
- Primary claims have authoritative support (gov/court/edu when available)
- No critical information gaps for practical guidance
- No unresolved contradictions
- Jurisdiction and dates are appropriate
Mark as INSUFFICIENT when:
- Missing critical legal requirements or procedures
- Only user-generated content for key claims
- Significant contradictions need resolution
- Wrong jurisdiction or outdated information
IMPORTANT for iteration {iteration_count + 1}:
- Be MORE LENIENT after multiple iterations
- Focus on whether user has enough info to take action
- Consider cumulative information across all iterations
SUGGESTED REFINEMENTS (if insufficient):
Provide 1-2 specific search queries that would help find the missing information.
These should be actual search queries that can be executed directly, NOT instructions.
Good examples:
- "California SB 365 employment arbitration 2025 full text"
- "Biden student loan forgiveness Supreme Court ruling 2024"
- "AB 465 arbitration agreement requirements California"
Bad examples (don’t do this):
- "Search for more information about SB 365"
- "Look for court rulings"
- "Find government sources"
Provide detailed reasoning for your judgment.
Appendix B Case Study: Executive Order Timeline Interpretation
This section presents a comparative analysis of a legal multiple-choice question where the L-MARS system with retrieval-augmented generation successfully identified the correct answer, while GPT-4o without external search failed.
Question ID 32: Executive Order Timeline
Question: According to Section 3 of the May 23, 2025 Executive Order
’Restoring Gold Standard Science,’ how soon must the OSTP Director
issue guidance for agencies on implementing ’Gold Standard Science’?
Answer Choices:
A: Within 30 days of the date of the order
B: Within 90 days of the date of the order
C: Within 6 months of the date of the order
D: Within 1 year of the date of the order
Correct Answer: A
L-MARS Simple Mode Response (Correct)
Selected Answer: A (Within 30 days)
Reasoning: The search results from library.washu.edu explicitly state
that the order mandated a timeline of 30 days for the Office of Science
and Technology Policy to issue guidance. This is the most direct evidence
relating to Section 3 of the Executive Order. The system cross-referenced
multiple authoritative sources to confirm this specific timeline requirement
rather than relying on general patterns or assumptions about typical
executive order timelines.
Search Sources Used: 5 Web Search sources
1. whitehouse.gov - "Restoring Gold Standard Science"
   Official White House publication of the Executive Order with full text
2. justice.gov - "New Executive Order on ’Gold Standard Science’: FOIA..."
   Department of Justice guidance on FOIA implications and compliance timelines
3. $h$s.gov - "Implementing Gold Standard Science"
   Department of Health and Human Services implementation guidelines
4. lawbc.com - "White House OSTP Issues Agency Guidance for Gold..."
   Legal analysis of OSTP guidance requirements and deadlines
5. library.washu.edu - "Federal Agencies Respond to ’Gold Standard Science..."
   Academic repository with detailed timeline analysis stating explicitly:
   "Section 3 mandates 30-day timeline for OSTP guidance issuance"
Key Evidence Retrieved:
- Executive Order issued: May 23, 2025
- OSTP guidance deadline: Within 30 days (by June 22, 2025)
- Agencies must implement principles of scientific integrity
- Guidance covers conflicts of interest in research and policy
Performance Metrics:
- Processing Time: 13.08 seconds
- Confidence Score: 0.9 (90%)
- U-Score (Uncertainty): 0.611 (lower is better)
- Citation Score: 1.0 (perfect source attribution)
- Jurisdiction Score: 1.0 (correctly identified federal scope)
- Decisiveness Score: 0.5 (moderate - provided clear guidance)
- Model: OpenAI GPT-4o (with RAG)
Result: Correct
GPT-4o Response (Incorrect)
Selected Answer: B (Within 90 days)
Reasoning: The question specifies the timing requirement for issuing
guidance by the OSTP Director according to Section 3 of the Executive
Order. Based on the context and typical executive orders that set
implementation timelines for guidance issuance, 90 days is commonly
allocated as a reasonable period for action. Therefore, the best
answer is B: Within 90 days of the date of the order.
Result: Incorrect
B.1 Analysis
This case demonstrates a critical advantage of retrieval-augmented generation in legal question answering:
B.1.1 Why GPT-4o Failed
  1. 1.
    Reliance on General Patterns: GPT-4o relied on typical patterns in executive orders, assuming 90 days as a “reasonable period” without access to the specific order text.
  2. 2.
    Lack of Specific Knowledge: The model’s training data cutoff prevented it from having specific information about an executive order in May 2025.
  3. 3.
    Statistical Bias: The model defaulted to a common timeline (90 days) that appears frequently in government documents, demonstrating the limitation of pattern-based reasoning without factual verification.
B.1.2 Why L-MARS Succeeded
  1. 1.
    Direct Evidence Retrieval: L-MARS retrieved the actual executive order text from authoritative sources, finding the explicit 30-day requirement in Section 3.
  2. 2.
    Source Authority: The system identified and prioritized information from library.washu.edu, an authoritative academic source with direct access to government documents.
  3. 3.
    Evidence-Based Reasoning: Rather than relying on typical patterns, L-MARS based its answer on explicit textual evidence from the retrieved documents.
B.2 Implications for Legal AI Systems
This case study highlights several important considerations for legal AI applications.
  • Temporal Sensitivity: Legal questions often require access to current and specific documents that may not be found in training data.
  • Precision Requirements: Legal deadlines and requirements demand exact information rather than reasonable approximations.
  • Source Verification: The ability to cite specific authoritative sources is crucial for the credibility of legal research.
  • RAG Advantage: Retrieval-augmented generation provides a significant advantage for questions requiring specific factual information, particularly for recent or specialized legal documents.
This example underscores that while large language models possess strong reasoning capabilities, augmenting them with retrieval mechanisms is essential for reliable legal question answering, especially when dealing with specific statutory requirements, deadlines, and recent legal developments.

### v3

@@
 Abstract
-We present L-MARS (Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search), a system that reduces hallucination and uncertainty in legal question answering through coordinated multi-agent reasoning and retrieval.
-Unlike single-pass RAG, L-MARS decomposes queries into subproblems, issues targeted searches across heterogeneous sources (Serper web, local RAG, CourtListener case law), and employs a Judge Agent to verify sufficiency, jurisdiction, and temporal validity before answer synthesis.
-This iterative reasoning–search–verification loop maintains coherence, filters noisy evidence, and grounds answers in authoritative law.
-We evaluated L-MARS on LegalSearchQA, a new benchmark of 200 up-to-date multiple choice legal questions in 2025.
-The results show that L-MARS substantially improves the accuracy of the factual data, reduces uncertainty, and achieves higher preference scores for the human and LLM judges.
-Our work demonstrates that multi-agent reasoning with agentic search is a scalable, reproducible blueprint for deploying LLMs in high-stakes domains requiring precise legal retrieval and deliberation.
-The code is available at: https://github.com/boqiny/L-MARS.
+We present L-MARS (Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search), a multi-agent retrieval framework for grounded legal question answering that decomposes queries into structured sub-problems, retrieves evidence via agentic web search, filters results through a verification agent, and synthesizes cited answers.
+Existing legal QA benchmarks test either closed-book reasoning or retrieval over fixed corpora, but neither captures scenarios requiring current legal information.
+We introduce LegalSearchQA, a 50-question benchmark across five legal domains whose answers depend on recent developments that post-date model training data.
+L-MARS achieves 96.0% accuracy in LegalSearchQA, a 38.0% improvement over zero-shot (58.0%), while chain-of-thought prompting degrades performance to 30.0%.
+On Bar Exam QA [28], a reasoning-focused benchmark of 594 bar examination questions, retrieval provides negligible gains (+0.7 percentage points), consistent with prior findings.
+These results show that agentic retrieval dramatically improves legal QA when tasks require up-to-date factual knowledge, but the benefit is benchmark-dependent, underscoring the need for retrieval-focused evaluation.
+Code and data: https://github.com/boqiny/L-MARS.
 1 Introduction
-Large language models (LLMs) such as GPT (gpt_4o_system_card, ), Claude (anthropic2024claude35sonnet, ), and Gemini (gemini2024arxiv, ) are increasingly applied to legal tasks such as case law retrieval (locke2022caselawretrievalproblems, ), statutory interpretation, and automated legal assistance(rincon2021legalclinic, ).
-Their ability to generate fluent context-sensitive responses makes them attractive for legal services, but direct application often results in hallucinations (ji2023survey, ; xu2025hallucinationinevitableinnatelimitation, ) or uncertainty (huang2024surveyuncertaintyestimationllms, )—confidently stated yet factually unsupported answers.
-In legal contexts, such errors carry significant real-world risk, as incorrect citations or outdated statutes can undermine credibility and thus negatively affect decision-making.
-Two primary strategies are commonly used to mitigate hallucinations in domain-specific LLMs.
-Domain-specific fine-tuning(chalkidis2020legalbert, ) can improve performance in legal queries, but is costly and fragile.
-Legal systems evolve rapidly: new regulations are introduced, existing laws are amended, and obsolete provisions are repealed.
-Maintaining accuracy would require frequent retraining, which makes this approach impractical for continuously updated domains.
-Retrieval-Augmented Generation (RAG) (lewis2020rag, ) avoids the retraining burden by retrieving relevant documents from an external corpus at inference time, but its effectiveness depends heavily on retrieval accuracy.
-If key legal evidence is missed, the model may still hallucinate or produce incomplete reasoning.
-To address these challenges, we introduce L-MARS, a system that integrates diverse retrieval sources with structured outputs and multi-turn model reasoning.
-L-MARS combines online search via the Serper API (serperapi, ) for up-to-date web information, offline retrieval using RAG over a curated local legal database, and case law retrieval through the CourtListener API (courtlistenerapi, ) for authoritative legal opinions.
-Unlike prior single-turn pipelines, L-MARS employs a multi-agent workflow that iteratively decomposes queries, executes targeted retrieval, and performs sufficiency checks before synthesizing final answers.
+Large language models (LLMs) such as GPT [1], Claude [3], and Gemini [4] are increasingly applied to legal tasks including case law retrieval [26], statutory interpretation, and automated legal assistance [2].
+However, direct application often produces hallucinations [5, 27] where the model states confidently incorrect legal facts.
+In legal contexts such errors carry real-world risk, as wrong citations or outdated statutes can undermine credibility and affect decision-making.
+Domain-specific fine-tuning [6] can help but requires frequent retraining as laws change, making it impractical for continuously evolving domains.
+Retrieval-Augmented Generation (RAG) [7] avoids retraining by fetching relevant documents at inference time, but its effectiveness depends on retrieval quality.
+Recent work has shown that on reasoning-intensive legal benchmarks, standard retrieval pipelines barely improve downstream QA accuracy.
+On Bar Exam QA, even the best traditional retriever improves GPT-4o-mini accuracy by only 0.5 percentage points over zero-shot [28].
+We introduce L-MARS, a multi-agent retrieval framework that addresses these challenges through structured agent orchestration.
+L-MARS combines online search via the Serper API [8], offline retrieval using BM25 over local documents, and case law retrieval through the CourtListener API [9].
+Unlike naive retrieve-and-generate pipelines, L-MARS employs a multi-agent workflow that decomposes queries into structured search intents, executes targeted retrieval, filters evidence through a Judge Agent, and synthesizes grounded answers.
 Figure 1: Comparison of the two L-MARS operating modes.
@@
 L-MARS operates in two modes.
-In the Simple Mode, a single-turn pipeline retrieves relevant legal documents and uses a model with reasoning capacity to integrate them into the final answer.
-In the Multi-Turn Mode, a Query Agent refines the user query through clarifying sub-questions, a Judge Agent evaluates the sufficiency and relevance of retrieved evidence, and a Summary Agent synthesizes the final answer.
-This iterative loop enables the system to adaptively search, verify, and refine until an evidence sufficiency threshold is met.
-By combining reasoning with agentic search, L-MARS achieves higher factual accuracy and significantly reduces uncertainty compared to the baseline of the GPT, Claude, and Gemini models.
-The framework also provides a flexible architecture that can switch between low-latency simple mode and high-accuracy multi-turn mode, offering a reproducible blueprint for high-stakes domains such as law.
-Table 1: Examples of Model-Expressed Uncertainty in Legal QA Responses
-Model Uncertainty Expressions in Pure LLM Responses for Legal Questions
-Regulations may differ across various cities and provinces in China.
-As of my last update in October 2023, public protest in China is heavily regulated…
-Legal interpretations and implementations may vary between EU Member States.
-This information is based on the legal framework as of 2023 and may be subject to change.
-It is important to note that regulations are subject to change, especially with new laws or court rulings.
+In Simple Mode, a single-pass pipeline decomposes the query, retrieves and filters evidence, and synthesizes a grounded answer.
+In Multi-Turn Mode, an iterative loop enables the system to adaptively search, verify, and refine until an evidence sufficiency threshold is met.
+In this work we evaluate Simple Mode to isolate the contribution of agent orchestration from iterative refinement.
+In legal practice, outdated information can lead to materially incorrect advice.
+Our results suggest that retrieval is essential for legal QA tasks that depend on current facts, and may be important for reliable deployment in domains where knowledge changes rapidly.
+Our contributions are threefold.
+First, we introduce LegalSearchQA, a benchmark for legal QA requiring post-training, time-sensitive knowledge.
+Second, we present L-MARS, a multi-agent retrieval framework for grounded legal question answering.
+Third, we show empirically that retrieval yields large gains only in retrieval-dependent settings, while chain-of-thought reasoning can degrade performance when parametric knowledge is outdated.
+All prompts, evaluation scripts, and raw outputs are released to ensure full reproducibility.
 2 Related Work
-Agentic RAG and Multi-Agent Framework
-RAG grounds the model output in external evidence to reduce hallucination.
-Moving beyond single-turn pipelines, interleaving reasoning with retrieval (yao2023reactsynergizingreasoningacting, ; khattab2023, ) improves factual grounding and sample efficiency.
-Multi-agent frameworks (wu2023autogen, ; langgraph2024, ) operationalize role specialization (planner, retriever, verifier) and conversational coordination, enabling iterative retrieval, critique, and revision.
-Recent agentic RAG variants further combine planning and verification with retrieval to maintain long-horizon reasoning(singh2025agenticRAG, ).
-Search Systems for LLMs
-Long-horizon tool-using research requires agents that actively plan queries, browse, and synthesize sources.
-Search-o1 and WebThinker (searcho1, ; li2025webthinkerempoweringlargereasoning, ) integrates agentic search into large reasoning models to trigger retrieval at uncertainty points and reason over documents before injection into the chain.
-In production settings, OpenAI’s Deep Research (openai2025deepresearch, ) exposes an agentic, web-scale research loop designed for deeper, broader investigations.
-These systems complement “think-then-answer” reasoning models (openai2024o1arxiv, ; openai2025o3systemcard, ) by coupling deliberation with targeted information acquisition.
-Reasoning Models and Techniques
-Reasoning-oriented prompting has shown strong gains: chain-of-thought (wei2023chainofthoughtpromptingelicitsreasoning, ) prompting encourages models to externalize intermediate steps, while self-consistency reduces variance by aggregating diverse traces (wang2022selfconsistency, ).
-Structured interaction methods, including debate-style critique and verifier-based confidence estimation, further improve factuality and calibration (kadavath2022know, ).
-More recently, large reasoning models such as OpenAI’s o1(openai2024o1arxiv, ), Qwen-QwQ(qwq-32b-preview, ), and DeepSeek-R1(deepseek-r1, ) extend these ideas by scaling the deliberation at the test time, explicitly demonstrating multi-step reasoning traces across domains such as mathematics, code, and science.
+Agentic RAG and Search Systems
+RAG grounds model output in external evidence to reduce hallucination [7], and interleaving reasoning with retrieval [30, 10] further improves factual grounding.
+Multi-agent frameworks [11, 12] enable role specialization and iterative critique, with recent agentic RAG variants combining planning, verification, and retrieval for long-horizon tasks [31, 33].
+On the search side, Search-o1 and WebThinker [13, 14] integrate agentic search into reasoning models, while OpenAI’s Deep Research [15] exposes a production-grade web-scale research loop.
+These systems complement reasoning models [16, 17] that scale deliberation at test time through chain-of-thought prompting [18], self-consistency [19], verifier-based calibration [32], and dedicated reasoning architectures [20, 21].
 Legal AI and Benchmarks
-Legal AI has progressed through a series of domain-specific benchmarks and corpora.
-LegalBench(guha2023legalbench, ) measures practical legal reasoning for diverse tasks ; LexGLUE(chalkidis2022lexglue, ) standardizes the evaluation across multi-task legal NLU; and Pile-of-Law(henderson2022pileoflaw, ) provides a large and responsibly curated pretraining corpus.
-More recent work has introduced reasoning and retrieval focused evaluations: Reasoning-Focused Legal Retrieval Benchmark(Zheng_2025, ) introduces Bar Exam QA and Housing Statute QA to capture realistic legal RAG settings, while LEXam (fan2025lexambenchmarkinglegalreasoning, ) benchmarks long-form, multi-step legal reasoning with law exam questions in English and German.
-Our work builds on this trajectory by combining retrieval with agentic multi-turn reasoning and domain-specific verification, complementing these benchmarks with a system architecture explicitly designed to reduce hallucination and uncertainty.
+Legal AI evaluation spans benchmarks for reasoning [24], multi-task NLU [22], Chinese legal knowledge [34], and pretraining data [23].
+More recent work targets retrieval and reasoning specifically: Bar Exam QA [28] captures realistic legal RAG settings where standard retrieval fails, and LEXam [29] benchmarks multi-step legal reasoning across languages.
+On the systems side, ChatLaw [35] explores multi-agent approaches for legal tasks.
+We note that many LegalBench [24] subtasks are now saturated, with GPT-4o-mini achieving 100% on the rule_qa, limiting their discriminative value for knowledge-intensive legal QA.
 3 Methodology
@@
 At each step, the system decides whether (i) expand $D$ by issuing new search queries, or (ii) terminate with a sufficient answer.
-Formally, the objective is to learn a mapping
+Formally, the system computes a mapping
 $(q,U,D)\;\mapsto\;(R,a),$
@@
 L-MARS is a multi-agent workflow implemented in LangGraph with structured output validation and tool use.
-The system employs a directed acyclic graph (DAG) architecture in which agents are represented as nodes and the control flow is represented as edges with conditional routing.
+The system employs a directed graph in which agents are represented as nodes and the control flow is represented as edges with conditional routing.
 The workflow maintains a centralized WorkflowState that tracks: (i) the evolving query representation, (ii) the accumulated search results with metadata, (iii) the iteration history and refinement notes, and (iv) the intermediate agent outputs.
@@
 Agents
-The Query agent parses the user’s question into a structured query result (issue type, key entities, time window, jurisdiction, and initial search intents).
+The Query agent parses the user’s question into a structured query result.
 The Search Agent executes tool calls to the retrieval backends and returns normalized SearchResult objects (title, url, snippet, optional full content, source type).
@@
 Online Search
-We adopt an agentic retrieval policy inspired by Search-o1(searcho1, ), interleaving targeted web queries with stepwise reasoning rather than performing a single problem-oriented fetch.
+We adopt an agentic retrieval policy inspired by Search-o1[13], interleaving targeted web queries with stepwise reasoning rather than performing a single problem-oriented fetch.
 Concretely, the Search Agent triggers retrieval when either (i) the Judge Agent flags a knowledge gap (e.g.
@@
 This preserves local evidence while discarding boilerplate and navigation chrome.
-The extractor also normalizes whitespace, strips punctuation, and truncates long pages to a hard cap to guaranty bounded latency.
+The extractor also normalizes whitespace, strips punctuation, and truncates long pages to a hard cap to guarantee bounded latency.
 Table 1: Comparison of Basic vs. Enhanced (Deep) Search on the query “Can I work remotely in the United States as an F1 student as of 2025?”
@@
 We also maintain a local retrieval index over user-provided documents.
-Markdown files are segmented into overlapping windows of 500 characters with 100-character stride and indexed using the BM25 ranking function (robertson2009bm25, ).
+Markdown files are segmented into overlapping windows of 500 characters with 100-character stride and indexed using the BM25 ranking function [25].
 At query time, the retriever returns the top-$k$ segments (default $k=5$), enriched with document titles and section headers to preserve context.
@@
 A final Summary Agent composes the answer.
-Algorithm 1 L-MARS Multi-Turn Workflow
-1:Input: user query $q$, max iterations $M$
-[…]
-[…]
-4:Receive user responses $U$;  $Q\leftarrow\textsc{Incorporate}(Q,U)$
-[…]
-6:for $i=1$ to $M$ do
-7:  $S\leftarrow\textsc{GenSearchQueries}(Q,R,E)$
-8:  $R_{i}\leftarrow\textsc{DeepSearch}(S,k{=}3,\text{with\_content}{=}\text{True})$
-9:  $(\text{suff},\text{notes})\leftarrow\textsc{JudgeAgent}(R\cup R_{i},T{=}0)$ $\triangleright$ CoT, quality, date/jurisdiction, contradiction
-10:  $R\leftarrow R\cup R_{i}$;  $E\leftarrow E\cup\text{notes}$
-11:  if $\text{suff}=\textsc{Sufficient}$ then
-12:   break
-13:  else
-14:   $Q\leftarrow\textsc{Refine}(Q,\text{notes},U)$
-15:  end if
-16:end for
-[…]
-18:return $(a)$
-Both modes share the same retrieval backends and dual evaluation stack.
-The key distinction lies in the graph topology: Simple Mode follows a linear path START $\to$ Query $\to$ Search $\to$ Answer $\to$ END, while Multi-Turn Mode includes a conditional loop with the Judge Agent determining whether to continue retrieval or proceed to answer generation based on evidence sufficiency criteria.
 Judge Agent and Stopping Rule
-By default, the Judge Agent is instantiated with GPT-o3 for stronger reasoning ability, and runs with temperature =0=0 to reduce variance and improve reproducibility.
+The Judge Agent runs with temperature =0=0 to reduce variance and improve reproducibility.
 Its prompt elicits three outputs: (i) a chain-of-thought style rationale, (ii) a structured checklist decision on sufficiency, and (iii) missing-evidence directives that guide the next refinement queries.
@@
 In practice, we found the Judge Agent especially useful when search results included unofficial forums (e.g., Reddit, Quora) instead of government domains; in such cases it flagged insufficiency and steered subsequent searches toward more compelling, authoritative sources such as government websites and official regulations.
+We describe both operating modes and all three retrieval backends for completeness, as they form the full L-MARS system available in our codebase.
+In our evaluation we use only Simple Mode with Serper web search, since this minimal configuration already achieves 96% accuracy on LegalSearchQA.
+Local RAG indexing and CourtListener case law retrieval are available as optional backends for practitioners working with proprietary document collections or case-specific research, but are not exercised in our benchmark evaluation.
 4 Experiments
 4.1 Tasks and Datasets
-We introduce LegalSearchQA, a benchmark of 200 legal questions designed to target post-training knowledge gaps.
-Unlike prior benchmarks such as LegalBench or CUAD(hendrycks2021cuadexpertannotatednlpdataset, ), which evaluate reasoning over pre-provided context, UncertainLegalQA tests the end-to-end ability of a system to retrieve and reason over external legal sources—an essential capability for real-world deployment.
-To avoid reliance on static pre-training knowledge, all questions explicitly reference the legal status as of 2025, requiring retrieval of up-to-date authorities.
-The dataset spans multiple domains, with a strong emphasis on federal executive orders (60%), along with coverage of tax provisions (10%), immigration (10%), federal tax inflation updates (10%), housing reforms (5%) and technology law (5%).
-This distribution reflects a focus on recent federal regulatory actions while still incorporating broader legislative and policy areas.
-Jurisdictional coverage includes both US federal law and major state-level developments, emphasizing contexts of regulatory flux, pending rulemaking, and areas where conflicting authorities require active retrieval.
-The questions are further categorized into requirement checks (48%), permissibility assessments (12%), definitions and scope (10%), policy intent (8%), timeline queries (8%), institutional authority (6%), legal authority references (4%) and eligibility thresholds (4%).
-Figure 2: Composition of the LegalSearchQA dataset.
-4.2 Evaluation Metrics
-We employ a comprehensive evaluation framework that combines automated metrics with LLM-based judgment to capture both quantitative and qualitative aspects of LegalSearchQA.
-Accuracy
-measures the fraction of multiple choice questions answered correctly against expert-annotated ground truth, reflecting the model’s ability to select the legally correct option.
-U-Score
-We propose U-Score, a rule-based metric for the evaluation of legal QA systems based on uncertainty.
-It ranges from 0 to 1 (lower is better) and captures reliability across five complementary dimensions: (i) hedging cues: frequency of uncertainty markers such as “may” or “could”; (ii) temporal vagueness: imprecise or underspecified references to time (e.g., “recently” instead of a year); (iii) citation sufficiency and authority: rewarding the use of primary legal sources (statutes, regulations, case law) over secondary commentary; (iv) jurisdictional specificity: explicit recognition of the relevant legal jurisdiction (e.g., U.S.
-federal vs.
-California state law); and (v) decisiveness: whether the answer delivers a clear conclusion rather than remaining equivocal.
-These components are linearly aggregated with weights:
-$\textbf{U-Score}=0.25\,H+0.20\,T+0.25\,(1-C)+0.15\,(1-J)+0.15\,(1-D),$ (1)
-where $H,T,C,J,D\in[0,1]$ denote hedging, temporal vagueness, citation sufficiency, jurisdictional specificity, and decisiveness, respectively.
-LLM-as-Judge
-complements these metrics with qualitative ratings using GPT-o3, evaluating answers along four dimensions: factual accuracy (consistency with current law), evidence grounding (relevance and credibility of cited sources), clarity of reasoning (logical coherence and completeness) and uncertainty calibration (appropriate hedging in genuinely ambiguous areas).
-Each response is assigned an overall rating of low, moderate, or high, reflecting the holistic quality of legal reasoning.
-Together, this dual evaluation framework captures both surface-level correctness and deeper reasoning quality.
+LegalSearchQA
+Existing legal QA benchmarks test either closed-book reasoning on static exam questions or retrieval over fixed corpora.
+Neither reflects the real-world setting where a legal professional needs current information, such as recently enacted statutes, updated regulatory thresholds, or new court decisions, that was not available when the model was trained.
+To address this gap we introduce LegalSearchQA, a 50-question multiple-choice benchmark designed to evaluate retrieval-augmented legal QA systems on current legal knowledge.
+Each question has four answer choices and exactly one correct answer.
+Questions span five legal domains: Tax, Corporate & Financial Regulation (13), Labor & Employment (11), Immigration (9), Technology & Privacy (9), and Criminal, Drug & State Law (8).
+The benchmark targets three categories of legal knowledge: status questions about the current state of laws, threshold questions about specific numerical values such as dollar amounts or deadlines, and requirement questions about what rules mandate.
+All questions were authored by the paper’s authors and verified against primary legal authorities such as IRS.gov, USCIS.gov, SEC.gov, and Supreme Court opinions during March 2026.
+By construction, correct answers require information from 2024 to 2026 that is unlikely to appear in model training data, making web retrieval essential.
+To reduce author bias, each question is grounded in a verifiable primary legal source and the task is purely factual retrieval rather than subjective interpretation, limiting degrees of freedom in answer labeling.
+While the absolute accuracy of 96% is high, this reflects the benchmark’s design: questions are intentionally selected to isolate retrieval-dependent knowledge where access to current information provides a clear advantage.
+To further reduce potential bias, we release all questions with source citations, timestamps, and verification notes.
+Each question is constructed to have a single unambiguous factual answer grounded in primary legal sources (e.g., IRS, SEC, Federal Register).
+Because the task is factual rather than interpretive, correctness can be independently verified without subjective judgment.
+The complete dataset is available at https://huggingface.co/datasets/boqiny/LegalSearchQA.
+Bar Exam QA
+We additionally evaluate on Bar Exam QA [28], a legal QA benchmark drawn from the Reasoning-Focused Legal Retrieval Benchmark.
+The dataset consists of multistate bar examination questions that test legal reasoning across core areas of U.S.
+law.
+We use the publicly released Historical MBE subset of 594 questions with subject labels spanning six domains: Contracts (113), Torts (112), Constitutional Law (95), Evidence (93), Real Property (92), and Criminal Law (89).
+The authors of Bar Exam QA showed that standard retrieval pipelines achieve very low recall on this benchmark and their best retriever improves GPT-4o-mini accuracy by only 0.5 percentage points over zero-shot.
+We include this benchmark to confirm that retrieval provides negligible gains on reasoning-focused tasks.
+We note that the Historical MBE subset overlaps with the MMLU auxiliary training set [36] and may inflate zero-shot baselines through memorization; we mitigate this by focusing primary claims on LegalSearchQA, where memorization is not a factor.
+4.2 Baselines and Experimental Conditions
+We compare three conditions, all using GPT-4o-mini as the backbone model:
+Zero-Shot
+The model answers directly with no retrieval and no explicit reasoning instructions. The prompt presents the question and four choices, instructing the model to select a single answer letter.
+Chain-of-Thought (CoT)
+The model is prompted to reason step-by-step: identify the legal issue, recall the relevant rule, apply it to the facts, and eliminate incorrect options before selecting an answer.
+No external retrieval is provided.
+L-MARS (Simple Mode)
+The full L-MARS pipeline in single-pass mode: the Query Agent decomposes the question into structured search intents, the Search Agent retrieves from Serper with enhanced content extraction, and the Summary Agent synthesizes the final answer with citations.
+All agents use GPT-4o-mini with T=0T{=}0.
+This experimental ladder isolates specific contributions: Zero-Shot to CoT tests whether reasoning alone helps, and CoT to L-MARS tests whether agentic retrieval adds value beyond reasoning.
+We leave evaluation of the Multi-Turn Mode to future work, as it targets open-ended legal research questions where a single retrieval pass may not surface sufficient evidence, a setting not captured by our current multiple-choice benchmarks.
+While L-MARS includes multiple components, our evaluation isolates the contribution of retrieval itself by comparing zero-shot against the full pipeline.
+Since Simple Mode excludes iterative refinement and Judge Agent loops, the observed improvements can be attributed primarily to structured web retrieval rather than multi-turn reasoning.
 4.3 Implementation Details
-All systems were evaluated on the 200-question LegalSearchQA benchmark using a standardized pipeline.
-The baseline LLMs (GPT-4o, Claude-4-Sonnet, Gemini-2.5-Flash) were run with temperature =0=0 and a 2048 token limit, using structured output or JSON mode to ensure consistency.
-Each model produced both a multiple choice answer and a supporting explanation, which were aggregated into structured JSON for evaluation.
-Performance was assessed along three complementary axes: accuracy, U-Score, and LLM-as-Judge.
-For reporting, accuracy and U-Score are averaged across all questions, while LLM-as-Judge ratings are determined by majority vote among per-question judgments.
-The complete evaluation pipeline, including prompts and scripts, is provided in the eval/ directory of our repository.
-4.4 Results
-Main Results on LegalSearchQA
-Table 3 reports aggregate results across three conditions: pure LLM inference, L-MARS in simple mode, and L-MARS in multi-turn mode.
-Our results highlight three key findings.
-First, both L-MARS variants substantially outperform pure LLM inference in accuracy (up to 98% vs.
-86–89%).
-Second, the U-Score decreases sharply under L-MARS (from 0.55–0.62 to 0.39–0.42), showing a marked reduction in hedging, vagueness, and unsupported conclusions.
-Third, LLM-as-Judge ratings consistently rank L-MARS outputs higher, with the multi-turn variant producing more thorough and contextually grounded answers.
-Finally, we note a trade-off in response time: while baseline LLMs answer within 1-4 seconds, L-MARS incurs higher latency (13.6s for simple, 55.7s for multi-turn), reflecting the added retrieval and reasoning steps.
-The visualization of the results is shown in Figure 3, and a detailed case study can be found in Appendix B.
-Table 3: Average performance of baseline LLMs and our proposed L-MARS framework on the LegalSearchQA benchmark.
-Method Accuracy U-Score LLM-as-Judge Response Time (s)
-GPT-4o 0.89 0.55 Moderate 1.69
-Claude-4-Sonnet 0.88 0.62 Moderate 3.84
-Gemini-2.5-Flash 0.86 0.58 Moderate 1.87
-L-MARS (simple, ours) 0.96 0.42 High 13.62
-L-MARS (multi-turn, ours) 0.98 0.39 High 55.67
+All conditions were evaluated using a standardized pipeline.
+For LegalSearchQA, we evaluated all 50 questions under each of the three conditions.
+For Bar Exam QA, we evaluated the full 594-question subset under zero-shot and L-MARS conditions.
+For the zero-shot and CoT conditions, prompts instruct the model to select a single answer letter (A–D).
+For L-MARS, all agents use GPT-4o-mini with T=0T{=}0; CourtListener integration was disabled since neither benchmark requires specific case citations.
+On Bar Exam QA, we additionally tested L-MARS with the Tavily search backend as an alternative to Serper.
+The complete evaluation pipeline, including prompts, scripts, and all results, is available at https://github.com/boqiny/L-MARS.
+5 Results
+5.1 LegalSearchQA
+Table 2 reports accuracy across three conditions on LegalSearchQA.
+L-MARS achieves 96.0% accuracy, improving over zero-shot by 38 percentage points while maintaining moderate latency at 2.4s per question.
+Chain-of-thought prompting degrades performance from 58.0% to 30.0%, falling below the 25% random baseline for four-choice questions.
+Without access to current information, CoT causes the model to construct plausible but outdated rationales and arrive at wrong answers with high confidence.
+We term this confident confabulation: the model recalls stale figures from prior tax years or superseded regulations, then builds step-by-step justifications around them.
+L-MARS rescued 19 questions that zero-shot answered incorrectly while regressing on none.
+We use a standard legal reasoning CoT prompt following the issue-rule-application-conclusion (IRAC) structure.
+We attribute the degradation to reliance on outdated parametric knowledge rather than prompt design, as the model produces internally coherent reasoning from stale premises across all five domains.
+We verified consistent degradation patterns with temperature T=0T{=}0 across all five domains, suggesting the effect is robust rather than an artifact of a single prompt configuration.
+Table 2: Accuracy on LegalSearchQA (50 questions). All conditions use GPT-4o-mini.
+Condition Retrieval Accuracy (%) Correct/Total Avg. Latency (s)
+Zero-Shot None 58.0 29/50 0.7
+Chain-of-Thought None 30.0 15/50 6.8
+L-MARS (Simple) Serper + agents 96.0 48/50 2.4
+Figure 2 and Table 3 break down accuracy by legal domain.
+The largest gains appear in Tax, Corporate & Financial Regulation, where L-MARS improves over zero-shot by 61.5 percentage points.
+Questions in this domain ask for specific dollar amounts, filing deadlines, and regulatory thresholds that changed in 2025 and 2026.
+Zero-shot accuracy here is only 30.8%, reflecting the model’s inability to access post-training information; L-MARS reaches 92.3% by retrieving current IRS and SEC sources.
+The smallest gains appear in Criminal, Drug & State Law, where zero-shot already achieves 87.5% because many questions concern well-publicized events that likely entered the model’s training data through news coverage.
+Notably, CoT falls below the random baseline in three of five domains, confirming that reasoning from stale knowledge is worse than no reasoning at all when current facts are required.
+Figure 2: Per-domain accuracy on LegalSearchQA. L-MARS achieves near-perfect accuracy across all domains. CoT falls below the 25% random baseline in three of five domains.
+Table 3: Per-domain accuracy (%) on LegalSearchQA. $\Delta$ is L-MARS gain over zero-shot.
+Domain N Zero-Shot CoT L-MARS $\Delta$
+Tax, Corporate & Financial 13 30.8 15.4 92.3 +61.5
+Labor & Employment 11 63.6 18.2 100.0 +36.4
+Technology & Privacy 9 55.6 44.4 100.0 +44.4
+Immigration 9 66.7 22.2 88.9 +22.2
+Criminal, Drug & State Law 8 87.5 62.5 100.0 +12.5
+Overall 50 58.0 30.0 96.0 +38.0
+5.2 Bar Exam QA
+Figure 3 visualizes the contrast between the two benchmarks.
+On Bar Exam QA (Table 4), L-MARS provides a negligible 0.7 point gain with Serper and a slight 0.8 point degradation with Tavily.
+Bar Exam QA tests legal reasoning over hypothetical fact patterns whose answers follow from general legal principles already in the model’s parameters, so web retrieval adds noise rather than signal.
+The gap between a 38 point gain on LegalSearchQA and a 0.7 point gain on Bar Exam QA confirms that the value of agentic retrieval depends on whether the task requires current information or abstract reasoning.
+Figure 3: Benchmark comparison. On retrieval-dependent LegalSearchQA, L-MARS gains 38 percentage points over zero-shot. On reasoning-focused Bar Exam QA, retrieval provides negligible improvement.
+Table 4: Accuracy on Bar Exam QA (594 questions). All conditions use GPT-4o-mini.
+Condition Acc. (%) Correct
+Zero-Shot 55.2 328/594
+L-MARS (Serper) 55.9 332/594
+L-MARS (Tavily) 54.4 323/594
+Prior work [28]:
+GPT-4o-mini 49.7
++ Best retriever 50.2
+Table 5: Per-subject accuracy (%) on Bar Exam QA.
+Subject N ZS L-MARS
+Const. Law 95 81.1 76.8
+Contracts 113 46.9 52.2
+Criminal 89 43.8 41.6
+Evidence 93 54.8 61.3
+Real Prop. 92 47.8 50.0
+Torts 112 57.1 53.6
+Table 5 shows per-subject results.
+L-MARS with Serper yields modest gains on Contracts (+5.3 points) and Evidence (+6.5 points) but slight regressions on Constitutional Law and Torts.
+No subject shows a large or consistent benefit from retrieval, reinforcing the conclusion that this benchmark primarily tests reasoning ability rather than information access.
 5.3 Error Analysis
-To better understand the limitations of the model, we conducted a qualitative error analysis of incorrect predictions.
-We observed two recurring patterns.
-First, models sometimes default to heuristic reasoning, such as assuming generic timelines or standard legal procedures rather than following the specific statutory text.
-Second, models occasionally overgeneralized from frequent legal phrases, leading to misclassification when a question involved narrow carving-outs or exceptions.
-These errors highlight two key challenges: reliance on surface-level priors instead of precise retrieval, and difficulty distinguishing closely related legal categories.
-Both issues underscore the importance of retrieval grounding and calibrated reasoning in complex legal QA.
-Figure 3: Average performance of baseline LLMs and our proposed L-MARS framework on the LegalSearchQA benchmark.
-Bars report Accuracy (blue) and $1-\text{U-Score}$ (purple), with response times shown alongside the $y$-axis labels.
-L-MARS substantially improves factual correctness and reduces uncertainty, though with higher latency compared to pure LLM inference.
-Human Expert Evaluation
-In addition, we conducted blind evaluations with graduate law students on a subset of 50 questions.
-Human preferences were closely aligned with LLM-as-Judge outcomes, with a 0.92 inter-annotator agreement rate.
-A recurring failure case was the Judge Agent’s tendency to over-reject partially relevant sources, occasionally triggering unnecessary search iterations.
-This suggests a direction for future work on trajectory-aware judge models that retain awareness of prior retrieval steps.
+L-MARS answered only two LegalSearchQA questions incorrectly.
+Q27 asked about H-1B weighted selection details involving a nuanced procedural rule about wage-level assignment for multiple registrations; the retrieved results did not surface the specific Federal Register provision.
+Q44 asked for the exact effective date of T+1 settlement; search results discussed the policy broadly but missed the precise date.
+Both errors share a pattern: the correct answer requires a specific fact buried in a longer document, and snippet-level retrieval missed the critical sentence.
+The CoT degradation deserves separate attention.
+On tax threshold questions, for instance, the model recalls outdated dollar amounts from prior years and builds multi-step justifications around them.
+This pattern is not unique to legal QA.
+In any domain where facts change faster than model training cycles, prompting the model to reason more carefully can amplify rather than correct errors, because the model reasons from wrong premises with greater conviction.
+This finding suggests caution when deploying CoT in fast-changing domains without concurrent retrieval.
+5.4 Ablation: Retrieval Depth
+[…]
 6 Conclusion
-We introduced L-MARS, a multi-agent workflow that integrates structured reasoning, agentic search, and sufficiency verification for legal question answering.
-By orchestrating specialized agents in iterative loops of query refinement, targeted retrieval, and evidence-based judgment, L-MARS significantly reduces hallucinations and uncertainty compared to pure LLM baselines.
-Our evaluation of the LegalSearchQA benchmark demonstrates consistent gains in factual accuracy, decisiveness, and grounding in authoritative sources.
-6 Limitations and Future Work
-While L-MARS improves reliability, its performance remains bounded by retrieval quality: if search engines or legal databases fail to return relevant authorities, downstream reasoning is impaired.
-The iterative multi-turn mode also introduces higher latency, which may be prohibitive in real-time legal assistance settings.
-Moreover, our evaluation focuses on US law; broader cross-jurisdictional and multilingual evaluations are needed to assess generalizability.
-Finally, the Judge Agent, though effective at flagging insufficiency, can be overly conservative, sometimes triggering unnecessary search iterations.
+We presented L-MARS, a multi-agent retrieval framework for legal QA, and LegalSearchQA, a 50-question benchmark requiring current legal knowledge that post-dates model training data.
+L-MARS achieves 96.0% on LegalSearchQA, a 38 percentage point gain over zero-shot, while CoT degrades accuracy to 30.0% through confident confabulation from stale knowledge.
+On reasoning-focused Bar Exam QA, retrieval provides negligible gains, confirming that the value of agentic retrieval is benchmark-dependent.
+These findings underscore the need for retrieval-focused legal benchmarks and suggest caution when deploying CoT in fast-changing domains without concurrent retrieval.
+Limitations
+LegalSearchQA contains 50 questions verified as of March 2026; the benchmark is sufficient for demonstrating large effect sizes but limits per-domain statistical power, and answers may become outdated as laws change.
+We evaluate only Simple Mode with web search; the Multi-Turn Mode, local RAG indexing, and CourtListener integration are available in the released codebase but not benchmarked here, as the simple configuration already demonstrates the core finding.
+Our evaluation covers U.S.
+law in English with a single backbone model, and performance remains bounded by search engine quality, as seen in the two errors where snippet-level retrieval missed critical details buried in longer documents.
 Appendix
 Appendix A Agent Instructions
-Query Agent Prompts
-# Follow-up Questions Generation Prompt
-You are a legal assistant helping users with their legal questions.
-User’s original question: {user_query}
-Generate 2-3 clarifying follow-up questions that would help you provide better
-legal assistance.
-Focus on:
-- Jurisdiction/location if not specified
-- Specific circumstances or context
-- Timeline or urgency
-- What type of help they need (information, next steps, etc.)
-Only ask questions that would significantly help in providing better assistance.
-# Search Query Generation Prompt
-Based on the user’s legal question and any additional context, generate 2-4 specific
-search queries.
-User Question: {user_query}
-Additional Context: {context}
-Generate queries for different sources:
-- case_law: For searching legal cases and precedents
-- web_search: For general legal information and recent updates
-- offline_rag: For searching local legal documents in the inputs folder
-Make queries specific and focused to get the most relevant results.
-# Query Analysis Prompt
-Analyze the user’s legal query and provide a detailed structured breakdown.
-User Query: {user_query}
-Additional Context: {context}
-Analyze and extract:
-1. Legal category/area (e.g., contract law, criminal law, family law, employment law)
-2. People involved with their roles (plaintiff, defendant, client, witness, etc.)
-3. Jurisdiction if mentioned or inferable
-4. Urgency level based on the nature of the query
-5. Specific legal areas involved
-6. Timeline or deadlines if mentioned
-7. Additional context about the situation
-Be thorough but accurate in your categorization.
-Summary Agent Prompt
-Create a comprehensive answer to the user’s legal question based on the search results.
-User’s Question: {user_query}
-Search Results:
-{results_content}
-Provide:
-1. A clear, comprehensive answer
-2. Key legal points and considerations
-3. Important disclaimers about legal advice
-Remember: This is informational only and not legal advice.
-Judge Agent Prompt
-You are a legal research judge evaluating search results.
-This is iteration {iteration_count + 1}.
-Original Question: {user_query}
-Conversation History:
-{conversation_context}
-Current Search Results ({len(search_results)} results):
-{results_summary}
-{prev_evaluations}
-EVALUATION PROTOCOL WITH CHAIN-OF-THOUGHT:
-1. REASONING (Chain of Thought):
-Think step by step about whether the search results answer the user’s specific question.
-Consider: What was asked? What information was provided? What is still missing?
-2. SOURCE QUALITY CHECK:
-Analyze source authority - are there sources from:
-- Government (.gov) sites?
-- Court decisions/legal databases?
-- Educational institutions (.edu)?
-- How many authoritative vs user-generated content sources?
-3. DATE CHECK:
-- Are the sources current and relevant to today’s date?
-- If there are older sources, do we also have recent confirmations?
-- Flag if critical information might be outdated
-4. JURISDICTION CHECK:
-- Does the jurisdiction of sources match the user’s location/scope?
-- For US federal vs state law, is the distinction clear?
-- User mentioned: {conversation_context if conversation_context
-else "No specific jurisdiction mentioned"}
-5. CONTRADICTION SCAN:
-- Do any sources contradict each other?
-- If yes, what specific elements conflict?
-- Do we need more specific queries to resolve conflicts?
-STOP RULE:
-Mark as SUFFICIENT when:
-- Primary claims have authoritative support (gov/court/edu when available)
-- No critical information gaps for practical guidance
-- No unresolved contradictions
-- Jurisdiction and dates are appropriate
-Mark as INSUFFICIENT when:
-- Missing critical legal requirements or procedures
-- Only user-generated content for key claims
-- Significant contradictions need resolution
-- Wrong jurisdiction or outdated information
-IMPORTANT for iteration {iteration_count + 1}:
-- Be MORE LENIENT after multiple iterations
-- Focus on whether user has enough info to take action
-- Consider cumulative information across all iterations
-SUGGESTED REFINEMENTS (if insufficient):
-Provide 1-2 specific search queries that would help find the missing information.
-These should be actual search queries that can be executed directly, NOT instructions.
-Good examples:
-- "California SB 365 employment arbitration 2025 full text"
-- "Biden student loan forgiveness Supreme Court ruling 2024"
-- "AB 465 arbitration agreement requirements California"
-Bad examples (don’t do this):
-- "Search for more information about SB 365"
-- "Look for court rulings"
-- "Find government sources"
-Provide detailed reasoning for your judgment.
-Appendix B Case Study: Executive Order Timeline Interpretation
-This section presents a comparative analysis of a legal multiple-choice question where the L-MARS system with retrieval-augmented generation successfully identified the correct answer, while GPT-4o without external search failed.
-Question ID 32: Executive Order Timeline
-Question: According to Section 3 of the May 23, 2025 Executive Order
-’Restoring Gold Standard Science,’ how soon must the OSTP Director
-issue guidance for agencies on implementing ’Gold Standard Science’?
-Answer Choices:
-A: Within 30 days of the date of the order
-B: Within 90 days of the date of the order
-C: Within 6 months of the date of the order
-D: Within 1 year of the date of the order
-Correct Answer: A
-L-MARS Simple Mode Response (Correct)
-Selected Answer: A (Within 30 days)
-Reasoning: The search results from library.washu.edu explicitly state
-that the order mandated a timeline of 30 days for the Office of Science
-and Technology Policy to issue guidance. This is the most direct evidence
-relating to Section 3 of the Executive Order. The system cross-referenced
-multiple authoritative sources to confirm this specific timeline requirement
-rather than relying on general patterns or assumptions about typical
-executive order timelines.
-Search Sources Used: 5 Web Search sources
-1. whitehouse.gov - "Restoring Gold Standard Science"
-   Official White House publication of the Executive Order with full text
-2. justice.gov - "New Executive Order on ’Gold Standard Science’: FOIA..."
-   Department of Justice guidance on FOIA implications and compliance timelines
-3. $h$s.gov - "Implementing Gold Standard Science"
-   Department of Health and Human Services implementation guidelines
-4. lawbc.com - "White House OSTP Issues Agency Guidance for Gold..."
-   Legal analysis of OSTP guidance requirements and deadlines
-5. library.washu.edu - "Federal Agencies Respond to ’Gold Standard Science..."
-   Academic repository with detailed timeline analysis stating explicitly:
-   "Section 3 mandates 30-day timeline for OSTP guidance issuance"
-Key Evidence Retrieved:
-- Executive Order issued: May 23, 2025
-- OSTP guidance deadline: Within 30 days (by June 22, 2025)
-- Agencies must implement principles of scientific integrity
-- Guidance covers conflicts of interest in research and policy
-Performance Metrics:
-- Processing Time: 13.08 seconds
-- Confidence Score: 0.9 (90%)
-- U-Score (Uncertainty): 0.611 (lower is better)
-- Citation Score: 1.0 (perfect source attribution)
-- Jurisdiction Score: 1.0 (correctly identified federal scope)
-- Decisiveness Score: 0.5 (moderate - provided clear guidance)
-- Model: OpenAI GPT-4o (with RAG)
-Result: Correct
-GPT-4o Response (Incorrect)
-Selected Answer: B (Within 90 days)
-Reasoning: The question specifies the timing requirement for issuing
-guidance by the OSTP Director according to Section 3 of the Executive
-Order. Based on the context and typical executive orders that set
-implementation timelines for guidance issuance, 90 days is commonly
-allocated as a reasonable period for action. Therefore, the best
-answer is B: Within 90 days of the date of the order.
-Result: Incorrect
+Query Agent
+The Query Agent is a stateless pass-through that wraps the user
+query into a structured QueryGeneration dataclass:
+    @dataclass
+    class QueryGeneration:
+        query: str
+        query_type: str = "web_search"
+        priority: str = "high"
+    class QueryAgent:
+        def build_query(self, user_query: str) -> QueryGeneration:
+            return QueryGeneration(query=user_query.strip())
+Summary Agent
+You are a bar exam expert. A user will provide retrieved legal
+sources and a multiple-choice question. Use the retrieved sources
+together with your legal knowledge to select the single best answer.
+Rules:
+1. Your response MUST start with exactly one letter on the first
+   line: A, B, C, or D.
+2. Follow with a concise 1-2 sentence rationale citing the
+   controlling legal rule.
+3. Bar exam questions test specific exceptions and nuances -- do
+   not just apply the most common general rule if the specific
+   facts call for an exception.
+Response format:
+[Letter]
+[Brief rationale]
+The user message is structured as:
+  Retrieved Evidence:
+  <concatenated search results with source, title, content>
+  Question:
+  <full MCQ with answer choices A-D>
+Judge Agent System Prompt
+You are a legal research quality judge for US bar exam questions.
+You will receive a bar exam multiple-choice question and a list of
+retrieved search result snippets. Your job is to decide whether the
+retrieved evidence is sufficient to reliably answer the question.
+"Sufficient" means: at least one result contains specific legal
+rules, statutes, case holdings, or doctrinal exceptions that
+directly address the core legal issue tested by the question -- not
+just a restatement of the facts.
+If the evidence is NOT sufficient, provide a better search query
+that targets primary legal sources (e.g., specific rule names,
+statute citations, or doctrine names) rather than generic question
+text.
+Respond with valid JSON only, no markdown fences:
+{"sufficient": true_or_false, "reason": "one sentence",
+ "next_query": "refined query (omit if sufficient)"}
+Pre-filtering: Before LLM evaluation, results matching any of
+these patterns are deterministically removed at zero cost:
+  HTTP error, 403/404, Connection error, Request timed out,
+  JavaScript is disabled, No search results found, etc.
+Appendix B Case Study: Post-Training Knowledge Gap
+We present two representative examples from LegalSearchQA where zero-shot and chain-of-thought both fail due to outdated parametric knowledge, while L-MARS retrieves the correct current information.
+Q1 (lsqa_001): Executive Order 14110 Status
+Question: What is the current status of Executive Order 14110, the
+Biden administration’s executive order on AI safety signed in
+October 2023?
+A. It remains in full effect and is being actively enforced
+B. It was revoked by President Trump on January 20, 2025
+C. It was partially amended but key provisions remain
+D. It expired automatically after 12 months
+Correct Answer: B
+Zero-Shot Response (Incorrect)
+Selected: A
+The model’s parametric knowledge contains EO 14110 as an active,
+enforceable order -- it has no information about the January 2025
+revocation, which post-dates its training data.
+Chain-of-Thought Response (Incorrect)
+Selected: A
+The model reasons step-by-step from stale knowledge: it recalls
+that EO 14110 was signed in October 2023, notes that executive
+orders do not expire automatically, and concludes it must still be
+in effect. The reasoning is internally consistent but factually
+wrong -- an example of "confident confabulation" from outdated
+parametric memory.
+L-MARS Response (Correct)
+Selected: B
+Executive Order 14110 has been revoked, as indicated by the sources
+stating that President Trump took action to revoke it during his
+final week in office. Therefore, it is no longer in effect.
+Retrieved sources included whitehouse.gov and Federal Register
+entries confirming revocation via EO 14148 on January 20, 2025.
+Q2 (lsqa_002): 2025 Standard Deduction
+Question: For tax year 2025, what is the standard deduction for a
+single filer under age 65, after the One Big Beautiful Bill Act
+amendments?
+A. \$14,600    B. \$15,000    C. \$15,350    D. \$15,750
+Correct Answer: D
+Zero-Shot: C (Incorrect)   CoT: C (Incorrect)   L-MARS: D (Correct)
+Zero-Shot selected C (\$15,350) -- the model recalls an intermediate
+figure from pre-OBBBA IRS projections but lacks knowledge of the
+legislative amendment.
+CoT selected C (\$15,350) -- step-by-step reasoning reinforces the
+wrong figure: the model recalls the 2024 standard deduction
+(\$14,600), applies an estimated inflation adjustment, and arrives
+at \$15,350. The reasoning is plausible but based on stale data.
+L-MARS selected D (\$15,750) -- retrieved current IRS sources
+confirming the One Big Beautiful Bill Act raised the 2025 standard
+deduction to \$15,750 for single filers.
 B.1 Analysis
-This case demonstrates a critical advantage of retrieval-augmented generation in legal question answering:
-B.1.1 Why GPT-4o Failed
-  1. 1.
-    Reliance on General Patterns: GPT-4o relied on typical patterns in executive orders, assuming 90 days as a “reasonable period” without access to the specific order text.
-  2. 2.
-    Lack of Specific Knowledge: The model’s training data cutoff prevented it from having specific information about an executive order in May 2025.
-  3. 3.
-    Statistical Bias: The model defaulted to a common timeline (90 days) that appears frequently in government documents, demonstrating the limitation of pattern-based reasoning without factual verification.
-B.1.2 Why L-MARS Succeeded
-  1. 1.
-    Direct Evidence Retrieval: L-MARS retrieved the actual executive order text from authoritative sources, finding the explicit 30-day requirement in Section 3.
-  2. 2.
-    Source Authority: The system identified and prioritized information from library.washu.edu, an authoritative academic source with direct access to government documents.
-  3. 3.
-    Evidence-Based Reasoning: Rather than relying on typical patterns, L-MARS based its answer on explicit textual evidence from the retrieved documents.
-B.2 Implications for Legal AI Systems
-This case study highlights several important considerations for legal AI applications.
-  • Temporal Sensitivity: Legal questions often require access to current and specific documents that may not be found in training data.
-  • Precision Requirements: Legal deadlines and requirements demand exact information rather than reasonable approximations.
-  • Source Verification: The ability to cite specific authoritative sources is crucial for the credibility of legal research.
-  • RAG Advantage: Retrieval-augmented generation provides a significant advantage for questions requiring specific factual information, particularly for recent or specialized legal documents.
-This example underscores that while large language models possess strong reasoning capabilities, augmenting them with retrieval mechanisms is essential for reliable legal question answering, especially when dealing with specific statutory requirements, deadlines, and recent legal developments.
+Both examples illustrate the same failure mode: the model’s parametric knowledge is confidently wrong about post-training developments.
+In Q1, a major policy reversal (EO revocation) is invisible to the model.
+In Q2, a legislative amendment changes a specific dollar figure that the model can only approximate from outdated data.
+Chain-of-thought prompting amplifies these failures rather than mitigating them.
+By encouraging the model to reason step-by-step from stale premises, CoT produces internally coherent but factually incorrect justifications.
+This “confident confabulation” effect explains why CoT accuracy (30.0%) falls below even zero-shot (58.0%) on LegalSearchQA.
+L-MARS resolves both cases by retrieving current authoritative sources (whitehouse.gov, IRS.gov) and grounding the answer in retrieved evidence rather than parametric memory.
+The system’s ability to access post-training information is the primary driver of its 38-percentage-point improvement over zero-shot on this benchmark.